# Eigenstate Thermalization Hypothesis (ETH) for off-diagonal matrix elements in integrable spin chains

### v1

@@ III.2 Statistics of matrix elements between eigenstates involving different macrostates
Let us now discuss off-diagonal matrix elements between eigenstates that are in different thermodynamic macrostates. First, exact results for hard-core bosons and numerical ones for the Lieb-Liniger gas (see [26]) suggest that off-diagonal matrix elements decay as
@@
where ${M_{ij}^{\prime}}^{\scriptscriptstyle{\mathcal{O}}}$ depends on the observable and on the two macrostates. Here we focus on the infinite-temperature and the zero-temperature (i.e., the ground state) macrostates. Indeed, we numerically observe that it is in general challenging to reach the asymptotic regime $L\to\infty$, where (3) is expected to hold. This could be attributed to the eigenstates-to-eigenstates fluctuations within the same macrostate at finite size $L$. These fluctuations give rise to “overlaps” between the two macrostates, which make challenging to observe the asymptotic scaling (24). It is natural to expect that the choice of the macrostates with $\beta=0$ and $\beta=\infty$ is expected to minimize such finite-size corrections.
In Fig. 7 we show the histograms of the off-diagonal matrix elements $M_{ij}=\ln(|\langle E_{i}|{\mathcal{O}}|E_{j}\rangle|^{2})$ for the one-spin operator $E_{i}^{(11)}$ (cf. (16)). As in Section III.1, we show results for $L=40,80,120,$ and $160$ and fixed particle density $M=L/4$. By comparing with Fig. 3 (a) we observe that the typical value of $M_{ij}\approx 10^{3}$ is much larger than for off-diagonal matrix elements between eigenstates in the same macrostate, which is compatible with the faster exponential decay Eq. 3. To extract the probability distribution function of $M_{ij}$, we fit the data to the Gumbel distribution (19). The fits are reported with the continuous lines. The agreement between the data and the fits is not as good as for matrix elements between eigenstates in the same macrostate, although this could be a finite $L$ effect. To test the expected scaling (3), in the inset of Fig. 7 we plot $\left|M_{ij}\right|/L^{2}$. The figure shows that all the data for $L\geq 80$ collapse on the same curve. In Fig. 8 we perform a similar analysis for the two-spin operator $E_{i}^{(11)}E_{i+1}^{(22)}$, showing data for $L=32,40,$ and $48$ and $M=L/4$. Again, the qualitative behavior of the data is similar as in Fig. 7, in agreement with Eq. 3.
@@ Appendix A Matrix elements of local operators in the XXZ chain
In this appendix, we review the exact result of Refs. [30, 31] for the matrix elements of the elementary operators (16) between energy eigenstates of the $XXZ$ spin-$1/2$ chain. The Hamiltonian of the model is
@@
where, as in the main text, $\vec{\sigma}_{i}$ are the Pauli matrices and the chain has length $L$. The parameter $\Delta$ is the anisotropy of the spin-spin interaction and, in the isotropic limit $\Delta=1$, the model (25) reduces to the $XXX$ spin chain (5) studied in the main text. The Hamiltonian (25) has a global $U(1)$ symmetry corresponding to the conservation of the total magnetization $S^{z}=1/2\sum_{i=1}^{L}\sigma_{i}^{z}$. As mentioned in the main text, at $\Delta=1$ this symmetry is enhanced to the full $SU(2)$ group of global spin rotations. In the following, in order to treat both the case $\Delta=1$ and $10<\Delta<$ 1 in an unified way, we define
$\phi(\lambda)$ =$=\begin{cases}\lambda,&\Delta=1,\\ \sinh{\lambda},&0<\Delta<1,\end{cases}$ (26)
$\eta$ =$=\begin{cases}-\mathrm{i},&\Delta=1,\\ -\operatorname{arccosh}{\Delta},&0<\Delta<1.\end{cases}$ (27)
@@
$\begin{split}T(\lambda;\{\xi_{i}\})&=R_{0N}(\lambda-\xi_{N})\ldots R_{01}(\lambda-\xi_{1})\\ &=\begin{pmatrix}A(\lambda;\{\xi_{i}\})&B(\lambda;\{\xi_{i}\})\\ C(\lambda;\{\xi_{i}\})&D(\lambda;\{\xi_{i}\})\end{pmatrix},\end{split}$ (29)
@@
The monodromy matrix (29) allows us to construct the Fock space of the model. Starting from the pseudovacuum state |$\ket{0}$, which for the $XXZ$ chain is the ferromagnetic state with all spins up $|\ket{0}=\ket{\uparrow\uparrow\uparrow\uparrow\ldots}$, the matrices $B(\mu)$ and $C(\mu)$ in Eq. 29 act on the pseudovacuum |$\ket{0}$ respectively as creation and annihilation operators of particles with rapidities $\mu_{i}$
|$\ket{\{\mu\}}=\prod_{k=1}^{M}B(\mu_{k})\ket{0},\quad\bra{\{\mu\}}=\bra{0}\prod_{k=1}^{M}C(\mu_{k})\,.$ (30)
We remark that not all particle states (30) are eigenstates of the Hamiltonian (25) and, as a consequence, they are not all orthogonal to each other, forming an overcomplete basis. To find the eigenstates of Eq. 25, we define the transfer matrix $\mathcal{T}$ as the trace over the internal space of the monodromy matrix (29) $\mathcal{T}(\lambda)=\operatorname{Tr}T(\lambda;\{\xi_{i}\})=A(\lambda;\{\xi_{i}\})+D(\lambda;\{\xi_{i}\})$. The transfer matrix evaluated at two generic rapidities $\lambda,\mu$ commute with each other $\left[\mathcal{T}(\lambda),\mathcal{T}(\mu)\right]=0$, meaning that they can be diagonalized simultaneously. Since $\lambda,\mu$ are arbitrary, this implies that we can diagonalize $\mathcal{T}(\lambda)$.
@@
Notice in particular that Eq. 32 vanishes when computed on the inhomogeneities, i.e., $d(\xi_{i})=0$. The results for the $XXZ$ chain are obtained by taking the homogeneous limit $\xi_{i}\to\eta/2$. Importantly, the homogeneous limit is plagued by fictitious singularities that have to be removed to recover physical results.
Moreover, as discussed already in Section II, the Bethe equations (31) admit solutions containing both real and complex rapidities. Solving the Bethe equations for generic complex rapidities is a daunting task. However, according to the string hypothesis [36], in the thermodynamic limit $L\to\infty$, the vast majority of complex solutions organizes in groups of $n$ rapidities known as Bethe string, which have identical real part and imaginary parts equally spaced by $\eta$. Physically, these strings states correspond to bound states of $n$ excitations moving collectively. In practice, at finite $L$ the string hypothesis is valid only approximately, and the rapidities in the same string exhibit deviations $\delta$ which are exponentially small in $L$. Assuming that the string hypothesis holds exactly even at finite (large) $L$ allows one to simplify considerably the solution of Eq. 31. By taking the limit of zero string deviations $\delta\to 0$ in (31), one obtains the Bethe-Gaudin-Takahashi equations (9) for the real parts of the strings. Again, the limit $\delta\to 0$ introduces additional fictitious singularities in the ABA formulas for the matrix elements.
@@ A.1 Scalar product of Bethe states
A crucial ingredient in the ABA formulas for the matrix elements of local operators is the so-called Slavnov formula for scalar products between Bethe states [38]. Let |$\ket{\{\lambda\}}$ be a Bethe state, with $M$ rapidities satisfying the Bethe equations (31), and let |$\ket{\{\mu\}}$ be a state of the form (30) with $M$ arbitrary parameters $\mu$. Clearly, the state |$\ket{\{\mu\}}$ for generic $\{\mu\}$ is not an eigenstate of the $XXZ$ chain. The two states are generically not orthogonal and their scalar product is given by the Slavnov formula [51, 38]
$\braket{\{\mu\}}{\{\lambda\}}=\frac{\det\mathbf{S}}{\prod_{i=1}^{M}\prod_{j<i}\phi(\mu_{i}-\mu_{j})\,\phi(\lambda_{j}-\lambda_{i})}\,,$ (33)
@@
If $\{\mu\}$ are also solutions of the Bethe equations, |$\ket{\{\mu\}}$ is an eigenstate of the $XXZ$ chain, and we can use the Bethe equation (31) to replace $d(\mu_{b})=\prod_{i\neq b}\phi(\mu_{i}-\mu_{b}+\eta)/\phi(\mu_{i}-\mu_{b}-\eta)$ in the Slavnov matrix (34). As a consequence now $R_{b}$ becomes
@@
While the previous expressions are always well defined for states with only real rapidities, for Bethe strings Eq. 39 vanishes. The rows of the Slavnov matrix (34) are then singular, although the determinant should remain finite. For simplicity, let us assume that $\{\mu\}$ contain only a single $n$-string, and let us rearrange $\{\mu\}$ such that the first $n$ rapidities are the ones forming the $n$-string. Following [32], we introduce string deviations $\delta$ for the first $n$ rapidities $\{\mu\}$. Then, we sum the first $n$ rows of the matrix (34). This operation remove the singularities without changing the determinant, allowing us to take the limit $\delta\to 0$. After these manipulations, the columns corresponding to the $n$ strings in the Slavnov matrix Eq. 34 are modified as
@@
The remaining columns of Eq. 34 are unchanged. Notice that in the case $n=1$, i.e., for a real rapidity, Eq. 40 reduces to Eq. 34, as expected. The procedure outlined above is straightforwardly generalized to the case in which $\{\mu\}$ contain an arbitrary number of strings [32] because different strings can be treated independently.
@@
$\braket{\{\lambda\}}{\{\lambda\}}=\frac{\det\mathbf{G}}{\prod_{i=1}^{M}\prod_{j\neq i}\phi(\lambda_{i}-\lambda_{j})}\,,$ (42)
@@
where the sum over $j>n$ in $\mathbf{G}^{\prime}_{an}$ is carried over all the rapidities that do not belong to the string, including possibly all other strings. Again we see that for $n=1$, Eq. 47 reduces to Eq. 43.
@@ A.2.1 General formulas for $\ell$-spin operators
One can exploit the fact that the Bethe state |$\ket{\{\lambda\}}$ is an eigenstate of the transfer matrix $\mathcal{T}$. Indeed, by using Eq. 49 and by applying the transfer matrix to the right eigenstate, we can write
$\bra{\{\mu\}}\prod_{i=1}^{\ell}E_{i}^{(\epsilon_{i},\epsilon^{\prime}_{i})}\ket{\{\lambda\}}=\Phi_{\ell}(\{\lambda\})\,\bra{\{\mu\}}\prod_{i=1}^{\ell}T_{\epsilon_{i},\epsilon_{i}^{\prime}}\ket{\{\lambda\}},$ (50)
@@
is the product of the eigenvalues of $\mathcal{T}(\xi_{j})$. To compute the matrix element we then apply the operators $A,B,C$ and $D$ in Eq. 29 on the state $\bra{\{\mu\}}$ | on the left. Recalling the definition (30) for the state $\bra{\{\mu\}}$ |, we observe that $C(\mu_{M+1})$ acts as a creation operator for an additional excitation with rapidity $\mu_{M+1}$. However, the action of $A,B$ and $D$ is more complicated and it is given as [30, 31, 24]
$\bra{0}\prod_{k=1}^{M}C(\mu_{k})\,A(\mu_{M+1})=\sum_{a^{\prime}=1}^{M+1}\frac{\prod_{k=1}^{M}\phi(\mu_{k}-\mu_{a^{\prime}}+\eta)}{\prod_{k=1,\,k\neq a^{\prime}}^{M+1}\phi(\mu_{k}-\mu_{a^{\prime}})}\left(\bra{0}\prod_{k=1,\,k\neq a^{\prime}}^{M+1}C(\mu_{k})\right),$ (52)
$\bra{0}\prod_{k=1}^{M}C(\mu_{k})\,D(\mu_{M+1})=\sum_{a=1}^{M+1}d(\mu_{a})\frac{\prod_{k=1}^{M}\phi(\mu_{k}-\mu_{a}-\eta)}{\prod_{k=1,\,k\neq a}^{M+1}\phi(\mu_{k}-\mu_{a})}\left(\bra{0}\prod_{k=1,\,k\neq a}^{M+1}C(\mu_{k})\right),$ (53)
$\begin{split}\bra{0}\prod_{k=1}^{M}C(\mu_{k})\,B(\mu_{M+1})=&\sum_{a=1}^{M}d(\mu_{a})\frac{\prod_{k=1}^{M}\phi(\mu_{k}-\mu_{a}-\eta)}{\prod_{k=1,\,k\neq a}^{M+1}\phi(\mu_{k}-\mu_{a})}\times\\ &\times\sum_{a^{\prime}=1,\,a^{\prime}\neq a}^{M+1}\frac{\prod_{k=1,\,k\neq a}^{M}\phi(\mu_{k}-\mu_{a^{\prime}}-\eta)}{\prod_{k=1,\,k\neq a,\,a^{\prime}}^{M+1}\phi(\mu_{k}-\mu_{a^{\prime}})}\left(\bra{0}\prod_{k=1,\,k\neq a,\,a^{\prime}}^{M+1}C(\mu_{k})\right).\end{split}$ (54)
In Eq. 54, we used the fact that in the case of interest $\mu_{M+1}=\xi$ and $d(\xi)=0$ to simplify the expression [30, 31, 24]. By recursively applying the outlined expressions, we obtain the action of a string of local operators on any Bethe state. To express the result in a closed form, we define the two set of indices $\alpha^{\pm}$ as [30, 31, 24]
@@
$\begin{split}1\leq a_{j}\leq M+j,\,a_{j}\in\mathbf{A}_{j}\,\text{ for }j\in\alpha^{-},\\ 1\leq a^{\prime}_{j}\leq M+j,\,a^{\prime}_{j}\in\mathbf{A}^{\prime}_{j}\,\text{ for }j\in\alpha^{+},\end{split}$ (56)
@@
$\begin{split}\mathbf{A}_{j}&=\big{\{}b\in[1,M+\ell]:b\neq a_{k},\,a_{k}^{\prime}\text{ with }k<j\big{\}},\\ \mathbf{A}_{j}^{\prime}&=\big{\{}b\in[1,M+\ell]:b\neq a_{k}\text{ with }k<j,\\ &\hskip 91.04872pt\text{ and }b\neq a_{k}^{\prime}\text{ with }k\leq j\big{\}}.\end{split}$ (57)
@@
$\bra{0}\prod_{k=1}^{M}C(\mu_{k})\,\prod_{j=1}^{\ell}T_{\epsilon_{j}\epsilon_{j^{\prime}}}(\mu_{M+j})$ (58)
$=$ $\sum_{\{a_{j},a_{j}^{\prime}\}}G_{\{a_{j},a_{j}^{\prime}\}}(\mu)\left(\bra{0}\prod_{b\in\mathbf{A}_{s+1}}C(\mu_{b})\right),$
@@
×$\times\prod_{j\in\alpha^{+}}\!\frac{\prod_{b=1,\,b\in\mathbf{A}^{\prime}_{j}}^{M+j-1}\phi(\mu_{b}-\mu_{a^{\prime}_{j}}+\eta)}{\prod_{b=1,\,b\in\mathbf{A}_{j+1}}^{M+j}\phi(\mu_{b}-\mu_{a^{\prime}_{j}})}.$

### v2

@@ III.2 Statistics of matrix elements between eigenstates involving different macrostates
Let us now discuss off-diagonal matrix elements between eigenstates that are in different thermodynamic macrostates. Notice that the results that we present here go beyond the standard ETH framework, which applies to matrix elements between eigenstates in the same microcanonical window. First, exact results for hard-core bosons and numerical ones for the Lieb-Liniger gas (see [26]) suggest that off-diagonal matrix elements decay as
@@
where ${M_{ij}^{\prime}}^{\scriptscriptstyle{\mathcal{O}}}$ depends on the observable and on the two macrostates. Here we focus on the infinite-temperature and the zero-temperature (i.e., the ground state) macrostates. Indeed, we numerically observe that it is in general challenging to reach the asymptotic regime $L\to\infty$, where (3) is expected to hold. This could be attributed to the eigenstates-to-eigenstates fluctuations within the same macrostate at finite size $L$. These fluctuations give rise to “overlaps” between the two macrostates, which make challenging to observe the asymptotic scaling (25). It is natural to expect that the choice of the macrostates with $\beta=0$ and $\beta=\infty$ minimizes such finite-size corrections. It is important to stress that, while the ground state is represented by a legit macrostate in thermodynamic Bethe Ansatz (TBA) language, it is a zero entropy one.
In Fig. 8 we show the histograms of the off-diagonal matrix elements $M_{ij}=\ln(|\langle E_{i}|{\mathcal{O}}|E_{j}\rangle|^{2})$ for the one-spin operator $E_{i}^{(11)}$ (cf. (16)). As in Section III.1, we show results for $L=40,80,120,$ and $160$ and fixed particle density $M=L/4$. By comparing with Fig.˜3 (a) we observe that the typical value of $M_{ij}\approx 10^{3}$ is much larger than for off-diagonal matrix elements between eigenstates in the same macrostate, which is compatible with the faster exponential decay Eq.˜3. To extract the probability distribution function of $M_{ij}$, we fit the data to the Gumbel distribution (19). The fits are reported with the continuous lines. The agreement between the data and the fits is not as good as for matrix elements between eigenstates in the same macrostate, although this could be a finite $L$ effect. To test the expected scaling (3), in the inset of Fig.˜8 we plot $\left|M_{ij}\right|/L^{2}$. The figure shows that all the data for $L\geq 80$ collapse on the same curve. In Fig.˜9 we perform a similar analysis for the two-spin operator $E_{i}^{(11)}E_{i+1}^{(22)}$, showing data for $L=32,40,$ and $48$ and $M=L/4$. Again, the qualitative behavior of the data is similar as in Fig. 8, in agreement with Eq.˜3.
@@ Appendix A Matrix elements of local operators in the XXZ chain
In this Appendix, we review the exact result of Refs. [30, 31] for the matrix elements of the elementary operators (16) between energy eigenstates of the XXZXXZ spin-$1/2$ chain. The Hamiltonian of the model is
@@
where, as in the main text, $\vec{\sigma}_{i}$ are the Pauli matrices and the chain has length $L$. The parameter $\Delta$ is the anisotropy of the spin-spin interaction and, in the isotropic limit $\Delta=1$, the model (26) reduces to the $XXX$ spin chain (5) studied in the main text. The Hamiltonian (26) has a global $U(1)$ symmetry corresponding to the conservation of the total magnetization $S^{z}=1/2\sum_{i=1}^{L}\sigma_{i}^{z}$. As mentioned in the main text, at $\Delta=1$ this symmetry is enhanced to the full $SU(2)$ group of global spin rotations. In the following, in order to treat both the case $\Delta=1$ and $0<\Delta<1$ in an unified way, we define
$\phi(\lambda)$ $=\begin{cases}\lambda,&\Delta=1,\\ \sinh{\lambda},&0<\Delta<1,\end{cases}$ (27)
$\eta$ $=\begin{cases}-\mathrm{i},&\Delta=1,\\ -\operatorname{arccosh}{\Delta},&0<\Delta<1.\end{cases}$ (28)
@@
[…]
@@
[…]
We remark that not all particle states (31) are eigenstates of the Hamiltonian (26) and, as a consequence, they are not all orthogonal to each other, forming an overcomplete basis. To find the eigenstates of Eq.˜26, we define the transfer matrix $\mathcal{T}$ as the trace over the internal space of the monodromy matrix (30) $\mathcal{T}(\lambda)=\operatorname{Tr}T(\lambda;\{\xi_{i}\})=A(\lambda;\{\xi_{i}\})+D(\lambda;\{\xi_{i}\})$. The transfer matrix evaluated at two generic rapidities $\lambda,\mu$ commute with each other $\left[\mathcal{T}(\lambda),\mathcal{T}(\mu)\right]=0$, meaning that they can be diagonalized simultaneously. Since $\lambda,\mu$ are arbitrary, this implies that we can diagonalize $\mathcal{T}(\lambda)$.
@@
Notice in particular that Eq.˜33 vanishes when computed on the inhomogeneities, i.e., $d(\xi_{i})=0$. The results for the $XXZ$ chain are obtained by taking the homogeneous limit $\xi_{i}\to\eta/2$. Importantly, the homogeneous limit is plagued by fictitious singularities that have to be removed to recover physical results.
Moreover, as discussed already in Section II, the Bethe equations (32) admit solutions containing both real and complex rapidities. Solving the Bethe equations for generic complex rapidities is a daunting task. However, according to the string hypothesis [36], in the thermodynamic limit $L\to\infty$, the vast majority of complex solutions organizes in groups of $n$ rapidities known as Bethe string, which have identical real part and imaginary parts equally spaced by $\eta$. Physically, these strings states correspond to bound states of $n$ excitations moving collectively. In practice, at finite $L$ the string hypothesis is valid only approximately, and the rapidities in the same string exhibit deviations $\delta$ which are exponentially small in $L$. Assuming that the string hypothesis holds exactly even at finite (large) $L$ allows one to simplify considerably the solution of Eq.˜32. By taking the limit of zero string deviations $\delta\to 0$ in (32), one obtains the Bethe-Gaudin-Takahashi equations (9) for the real parts of the strings. Again, the limit $\delta\to 0$ introduces additional fictitious singularities in the ABA formulas for the matrix elements.
@@ A.1 Scalar product of Bethe states
[…]
$\braket{\{\mu\}|\{\lambda\}}=\frac{\det\mathbf{S}}{\prod_{i=1}^{M}\prod_{j<i}\phi(\mu_{i}-\mu_{j})\,\phi(\lambda_{j}-\lambda_{i})}\,$, (34)
@@
[…]
@@
While the previous expressions are always well defined for states with only real rapidities, for Bethe strings Eq.˜40 vanishes. The rows of the Slavnov matrix (35) are then singular, although the determinant should remain finite. For simplicity, let us assume that $\{\mu\}$ contain only a single $n$-string, and let us rearrange $\{\mu\}$ such that the first $n$ rapidities are the ones forming the $n$-string. Following [32], we introduce string deviations $\delta$ for the first $n$ rapidities $\{\mu\}$. Then, we sum the first $n$ rows of the matrix (35). This operation removes the singularities without changing the determinant, allowing us to take the limit $\delta\to 0$. After these manipulations, the columns corresponding to the $n$ strings in the Slavnov matrix Eq.˜35 are modified as
@@
The remaining columns of Eq.˜35 are unchanged. Notice that in the case $n=1$, i.e., for a real rapidity, Eq.˜41 reduces to Eq.˜35, as expected. The procedure outlined above is straightforwardly generalized to the case in which $\{\mu\}$ contain an arbitrary number of strings [32] because different strings can be treated independently.
@@
$\braket{\{\lambda\}|\{\lambda\}}=\frac{\det\mathbf{G}}{\prod_{i=1}^{M}\prod_{j\neq i}\phi(\lambda_{i}-\lambda_{j})}\,$, (43)
@@
where the sum over $j>n$ in $\mathbf{G}^{\prime}_{an}$ is carried over all the rapidities that do not belong to the string, including possibly all other strings. Again we see that for $n=1$, Eq.˜48 reduces to Eq.˜44.
@@ A.2.1 General formulas for $\ell$-spin operators
[…]
@@
[…]
In Eq.˜55, we used the fact that in the case of interest $\mu_{M+1}=\xi$ and $d(\xi)=0$ to simplify the expression [30, 31, 24]. By recursively applying the outlined expressions, we obtain the action of a string of local operators on any Bethe state. To express the result in a closed form, we define the two set of indices $\alpha^{\pm}$ as [30, 31, 24]
@@
[…]
@@
[…]
@@
[…]
@@
$\times\prod_{j\in\alpha^{+}}\!\frac{\prod_{b=1,\,b\in\mathbf{A}^{\prime}_{j}}^{M+j-1}\phi(\mu_{b}-\mu_{a^{\prime}_{j}}+\eta)}{\prod_{b=1,\,b\in\mathbf{A}_{j+1}}^{M+j}\phi(\mu_{b}-\mu_{a^{\prime}_{j}})},$